# Semi-supervised Underwater Image Enhancement Using A Physics-Aware Triple-Stream Network

### v3

@@ Abstract
The images captured in the underwater environment play important roles in ocean exploration. However, these images normally suffer from different degradation, due to the wavelength-dependent light absorption and light scattering caused by the underwater transmission medium. In [1], the Image Formation Model (IFM) can be expressed as:
@@
Considering the degradation may interfere with downstream tasks, Underwater Image Enhancement (UIE) is key to underwater machine vision. Due to the complicated degradation process, this task is challenging. Many traditional methods [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] have been proposed, to address the challenge. These methods can be treated as the prior-based methods because they typically rely on a certain form of prior information. Some prior-based methods [3, 4, 5] aimed to estimate the parameters of the IFM. However, they normally struggle with adapting to diverse underwater scenes or estimating valid parameters when the scene violates the prior assumption [5]. On the other hand, the other prior-based methods [11, 12, 13, 14, 15] directly enhance the quality of underwater images without explicitly modeling the degradation process. Due to the lack of the knowledge of underwater scenes, they cannot effectively enhance the images captured in the complex scenes and often introduce artifacts.
@@
In contrast, deep learning-based UIE approaches [18, 19, 20, 21, 22, 23, 24, 25, 26] aim to directly learn from different types of degraded images. Therefore, the problem with the limited static priors can be alleviated. Ideally, a large number of real-world labeled training images should be utilized. However, it is difficult to collect such a large labeled underwater data set in practice. Alternatively, the best enhanced image can be selected as the reference (ground-truth) by observers from the images processed by a set of UIE algorithms [22, 27]. Nevertheless, this sort of data sets cannot represent diverse underwater scenes due to the limited size. In this case, the deep UIE model trained may overfit the training data.
@@
The synthetic images generated using the IFM [28, 18] or a Generative Adversarial Network (GAN) [29, 30, 31, 20, 21] were employed in order to address the problem with the lack of labeled data. However, the utilization of these images to simulate real-world underwater degradation processes is struggling. Transfer learning techniques [30, 19] were also applied for that purpose, which usually used generative models to map degraded images to the clearer images. But they are difficult to train and may produce unstable results. Recently, promising results were derived using zero-shot learning approaches [32, 33] by directly learning degradation parameters from unlabeled degraded images.
@@
To effectively model the degradation process, we propose a Physics-Aware Triple-Stream Underwater Image Enhancement Network, i.e., PATS-UIENet, by injecting the physical principle into the deep neural network. This network performs the UIE task by explicitly estimating the degradation parameters of a revised IFM [2] using a Direct Signal Transmission Estimation Steam (D-Stream), a Backscatter Signal Transmission Estimation Steam (B-Stream) and an Ambient Estimation Stream (A-Stream). These parameters are fed into the revised IFM and the result is the image that the PATS-UIENet enhances.
@@
We further adopt an IFM-inspired semi-supervised learning framework, including a bi-directional supervised scheme and an unsupervised scheme, to overcome the limitation of insufficient data. Compared with the unsupervised methods [32, 33], the framework can be more easily trained due to the facilitation from the supervised scheme. Since the unsupervised scheme uses additional unlabeled images while the bi-directional supervised scheme exploits labeled images more effectively, it can be better trained than the supervised approaches [18, 19, 22, 23, 24, 25, 26, 34]. As a result, the generalization of the model trained using our framework is stronger than that trained using a supervised or unsupervised method.
@@
We propose a physics-aware triple-stream network for the UIE task by explicitly estimating the degradation parameters of the revised IFM. Such an IFM-motivated network has not been explored for the UIE task before.
An IFM-inspired semi-supervised learning framework is adopted, to overcome the challenge of insufficient data. This framework exploits the advantages of both the supervised and unsupervised methods.
We conduct a series of comparative experiments on five real-world underwater testing sets along with sixteen baselines. These results provide the community with a set of benchmarks.
@@ II Related Work
II-A Prior-Based Methods
@@
II-B Learning-Based Methods
@@
In contrast to the lack of labeled underwater images, it is easier to collect unlabeled underwater images. Existing studies [19, 46] sourced real-world underwater images and categorized them according to the degree of degradation. Unsupervised learning approaches were developed on top of these data by learning the mapping from severely degraded images to slightly degraded images. However, these approaches were hard to train and tended to produce unstable results. Recently, a semi-supervised UIE method [47] was developed. Nevertheless, this method did not consider physical principles.
@@ III Methodology
Considering the importance of physical principles to the UIE task, we propose a Physics-Aware Triple-Stream Underwater Image Enhancement Network (PATS-UIENet). This network contains a Direct Signal Transmission Estimation Steam (D-Stream), a Backscatter Signal Transmission Estimation Steam (B-Stream) and an Ambient Light Estimation Stream (A-Stream), which are used to explicitly estimate the parameters of the revised IFM [2]. To address the challenge of the lack of labeled real-world underwater images, we further adopt an IFM-inspired semi-supervised learning framework, which consists of a bi-directional supervised scheme and an unsupervised scheme. Compared to the supervised or unsupervised method, the PATS-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model trained is thus stronger, due to the complementary action of the two schemes.
III-A Revised IFM
As pointed out by Akkaynak and Treibitz [2], the original IFM [1] ignored the dependencies of the backscatter coefficient on the ambient light and the optical property of water bodies. They assumed that the attenuation coefficients of the direct signal and the backscattered signal are the same. In this case, the original IFM [1] cannot describe the degradation process of underwater images adequately.
where and are two sets of parameters which affect the attenuation factors of the direct signal and the backscatter signal , respectively, represents the distance between the scene and the camera, is the reflectance spectrum of the object, is the ambient light at a certain distance, is the camera response function, is the beam scattering coefficient and is the beam attenuation coefficient. The above formula can be written in the more compact form as:
@@ III-B PATS-UIENet
Red Chanel Tuner (RCT). Inspired by the channel attention mechanism [49], we propose a compact RCT module in order to adaptively emphasize the information contained in the red channel. This module first uses a convolutional layer to extract basic features from the input image, and then use the Global Average Pooling (GAP) to obtain a global representation. A fully-connected layer and the Sigmoid activation function are further used to transform this representation to a tuning weight. The weight is finally used to scale the red channel of the image.
@@
D-Stream. According to Eq. (3), the transmission map varies at different pixel locations. In this case, the network should be able to generate fine-grained representations. Hence, we build the D-Stream using a CNN-based encoder-decoder network. The encoder comprises five consecutive blocks, denoted as to , and the Max-pooling operation is applied to each block. As a result, multi-scale feature maps are derived using the encoder. Given is used as the bottleneck, the decoder contains four blocks, denoted as to , symmetrically. Bilinear interpolation is used to upsample the feature maps within each block. Both the encoder and decoder blocks share the same structure, which consists of two -- units. Skip connection is used to pass the feature maps produced by an encoder block to the corresponding decoder block at the same level, which is useful for restoring the fine-grained spatial structure. The output of the last decoder block is fed into a convolutional layer. A set of transmission maps are produced in terms of different color channels.
B-Stream. Similarly, the B-stream is also designed as a CNN-based encoder-decoder network with the same architecture as the D-stream, motivated by Eq. (3). Unlike the D-stream, however, the output of the first encoder block in the B-stream undergoes the Patchify processing, which includes both the flattening and linear projection operations, and the result is then passed to the A-stream. Besides, a Residual Communication Module (RCM) connects the to in the B-stream with the to in the A-stream at the same level, respectively.
A-Stream. The ambient light describes the luminance of the scene. Since this parameter does not encode the information of a specific object, it can be treated as homogeneous. We are motivated to design the A-stream using a set of Transformer [50] blocks for the sake of leveraging the self-attention mechanism to extract the global information. Five blocks, denoted as to , are comprised of this stream. Each block contains a self-attention layer and a two-layer Multi-layer Perceptron (MLP) unit, in which two residual connections are used separately.
Since the input of the Transformer is a sequence of visual tokens, the feature maps of are processed using Patchify. Considering that the estimation of the ambient light does not require local characteristics and the quadratic complexity of self-attention is unavoidable, we downsample these maps by a factor of 16 to reduce the computational demand. The result is a set of tokens. A special token, namely, the Ambient token, is also adopted, to specifically learn the ambient light parameter. All the tokens are fed into . The Ambient token produced by the last block is sent to a linear layer. Finally, the estimation of the ambient light is achieved.
@@ III-C IFM-Inspired Semi-supervised Learning Framework
To overcome the challenge of insufficient real-world underwater images, we adopt a semi-supervised learning framework (see Fig. 3), inspired by the revised IFM [2] theory. This framework includes a bi-directional supervised scheme and an unsupervised scheme. The bi-directional scheme contains a forward-enhancement module, which learns from reference images and uses the estimated parameters of the revised IFM to perform UIE, and a backward-degradation module, which uses the estimated parameters to degrade the reference images to their real-world degraded counterparts. As a result, the bi-directional supervised scheme can learn a more accurate estimation of the parameters using limited real-world data than that learned using a supervised method with a large number of synthetic data. Also, the generalization of the model trained is stronger than that trained using a supervised method with limited real-world data. On the other hand, the unsupervised scheme exploits the unlabeled real-world underwater images that we collect. A second set of images with different degrees of degradation are generated from these images based on the revised IFM [2]. Two sets of parameters can be estimated using both sets of images, respectively. Each set is utilized as the reference for the other set because their IFMs are related.
@@
Unsupervised Learning. To leverage unlabeled underwater images, generative networks [30, 19] were used. However, they were difficult to train and produced unstable results. Since unsupervised learning methods [32, 33] are able to directly learn from degraded images without requiring reference images, we are motivated to adopt an unsupervised learning scheme on top of the revised IFM [2].
@@ IV-A Baselines
We compared the proposed method with four prior-based approaches which were developed on top of the IFM [1], including DCP [3], UDCP [4], GDCP [5] and DHL [10], four prior-based methods which did not consider the IFM, including Histogram Equalization (HE) [9], Retinex [13], MMLE [14] and HLRP [15], and eight deep learning approaches, including WaterNet [22], UColor [23], FUnIE-GAN [20], USUIR [24], PUIE-Net (MP) [25], PUIE-Net (MC) [25], U-Transformer [21] and URanker [26].
@@ IV-C Evaluation Metrics
For the labeled testing data, we used both Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [54] to assess the quality of an enhanced image with regard to the reference image. Regarding the unlabeled testing data, we used the Underwater Image Quality Measure (UIQM) [55] and Underwater Color Image Quality Evaluation (UCIQE) [56] metrics. When the SQUID [10] data set was tested, the Average Reproduction Angular Error () [10] was used to evaluate the quality of color restoration. Since transmission maps are related to the scene distance, we used the Pearson Correlation Coefficient (PCC) calculated between the estimated transmission maps and the ground-truth depth map to assess the performance of transmission estimation [10].
@@ IV-D Implementation Details
We implemented the proposed method using Pytorch and conducted experiments on Ubuntu 20.04 with a GeForce RTX 3090 graphics processing unit. During the training process, the images were resized to the resolution of pixels. The number of filters in the RCT was set to 16. For the to in the D-stream and B-stream, the numbers of filters were set to 64, 128, 256, 512, 512, 256, 128, 64 and 64 in turn. Regarding the A-stream, the dimension of each token was 384 and the number of attention heads in the MHSA was 6. We trained the PATS-UIENet using the AdamW [57] optimizer. The learning rate and batch size were set to 1e-4 and 12, respectively. Four weighting factors, including and , were set to 0.001, 0.005, 10 and 0.001, respectively. We first trained the PATS-UIENet using the bi-directional supervised scheme for 3,000 iterations as a warm-up stage. Then the semi-supervised scheme was used to train it for 150,000 iterations. When the semi-supervised learning process was carried out, in particular, the unsupervised scheme was performed for 30 iterations after the bi-directional supervised scheme was conducted for every 120 iterations.

### v4

@@ Abstract
The images captured in the underwater environment play important roles in ocean exploration. However, these images normally suffer from different degradation, due to the wavelength-dependent light absorption and light scattering caused by the underwater transmission medium. According to the Image Formation Model (IFM) [1], the process of image degradation can be formulated as follows:
@@
Considering that the degradation may interfere with downstream tasks, the design of effective UIE methods is critical yet challenging. Existing methods can be divided into two categories, i.e., traditional prior-based methods and deep learning-based methods. Traditional prior-based methods [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] typically rely on a certain form of prior information. Some prior-based methods [3, 4, 5] aim at estimating the parameters of the IFM. However, they normally struggle with adapting to diverse underwater scenes or estimating valid parameters when the scene violates the prior assumption [5]. On the other hand, the other prior-based methods [11, 12, 13, 14, 15] directly enhance the quality of underwater images without explicitly modeling the degradation process. Due to the lack of the knowledge of underwater scenes and the limited adaptability and robustness, they cannot effectively enhance the images captured in the complex scenes and often introduce artifacts.
@@
In contrast, deep learning-based UIE approaches [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] aim to directly learn from different types of degraded images. Therefore, the problem with the limited static priors can be alleviated. Ideally, a large number of real-world labeled training images should be utilized. However, it is difficult to collect such a large labeled underwater data set in practice. Alternatively, the best enhanced image can be selected as the reference (ground-truth) by observers from the images processed by a set of UIE algorithms [22, 31]. Nevertheless, this sort of data sets cannot represent diverse underwater scenes due to the limited size. In this case, the deep UIE model trained may overfit the training data.
@@
The synthetic images generated using the IFM [32, 18] or a Generative Adversarial Network (GAN) [33, 34, 35, 20] were used to address the problem with the lack of labeled data. However, the utilization of these images to simulate real-world underwater degradation processes is struggling. Transfer learning techniques [34, 19] were also applied for that purpose, which usually used generative models to map degraded images to the clearer images. But they are difficult to train and may produce unstable results. Although unsupervised learning methods [36, 37] alleviate the reliance on the paired data by directly learning degradation parameters from unlabeled degraded images, they often suffer from slow inference speed and are prone to color distortion. Recently, semi-supervised learning methods [38, 28, 29, 30], which combine the strengths of both the supervised and unsupervised approaches, have achieved promising results by leveraging the labeled data along with the unlabeled data during the training process.
Despite the progress has been made by the existing methods, two challenges remain. First, existing deep learning-based methods lack effective physics-aware modeling, which limits their interpretability and effectiveness in modeling the degradation process. Second, a supervised or unsupervised approach normally struggles with the insufficient or noisy data due to the complexity of real underwater degradation, leading to overfitting or weak generalization.
@@
To address these challenges, we propose a Physics-Aware Triple-Stream Underwater Image Enhancement Network, i.e., PATS-UIENet, which explicitly incorporates the revised IFM [2] into the deep neural network that we deliberately design. Unlike existing physics-based methods [3, 4, 5], our network explicitly estimates the parameters of a physical model using three streams, including a Direct Signal Transmission Estimation Stream (D-Stream), a Backscatter Signal Transmission Estimation Stream (B-Stream) and an Ambient Estimation Stream (A-Stream). Due to the explicit modeling of the physical degradation process, our network is superior to existing UIE methods by exploiting the strengths of both the theoretical interpretability of physics-based approaches and the powerful feature representation ability of deep neural networks.
@@
We further introduce an IFM-inspired semi-supervised learning framework, including a bi-directional supervised scheme and an unsupervised scheme, for the purpose of overcoming the limitation of insufficient data. The supervised scheme effectively utilizes the limited labeled data by providing explicit degradation guidance, while the unsupervised scheme exploits the abundant unlabeled data to improve generalization. Therefore, our framework can be better trained than the existing supervised approaches [18, 19, 22, 23, 24, 25, 26, 27], unsupervised methods [36, 37] and semi-supervised approaches [38, 28, 29, 30] using the same amount of data. As a result, the generalization of the model trained using our framework is stronger than that trained using existing supervised, unsupervised and semi-supervised approaches.
@@
We introduce a Physics-Aware Triple-Stream Underwater Image Enhancement Network, referred to as PATS-UIENet, which can explicitly estimate the degradation parameters of the revised IFM. This network differentiates itself from existing IFM-based methods by integrating physical parameter estimation into a deliberately designed deep neural network. Such an IFM-motivated network has not been explored for the UIE task before.
We propose an IFM-inspired semi-supervised learning framework, which addresses the issues of data insufficiency and training instability by exploiting the merits of both the supervised and unsupervised learning methods.
We conduct a series of comparative experiments on six underwater testing sets along with sixteen baselines. The results not only validate the effectiveness of our method but also provide benchmarks for future research.
@@ II Related Work
II-A Revised Image Formation Model
As pointed out by Akkaynak and Treibitz [2], the original IFM [1] ignored the dependencies of the backscatter coefficient on the ambient light and the optical property of water bodies. They assumed that the attenuation coefficients of the direct signal and the backscattered signal are the same. In this case, the original IFM [1] cannot adequately describe the degradation process of underwater images. It has been demonstrated that the attenuation factors of the direct signal and the backscatter signal are different while the attenuation factor of the backscatter signal is affected more severely by the ambient light [2].
where and are two sets of parameters which affect the attenuation factors of the direct signal and the backscatter signal , respectively, represents the distance between the scene and the camera, is the reflectance spectrum of the object, is the ambient light at a certain distance, is the camera response function, and and are the beam scattering and attenuation coefficients, respectively.
II-B Prior-Based Methods
@@
II-C Learning-Based Methods
@@
In contrast to the lack of labeled underwater images, it is easier to collect unlabeled underwater images. Existing studies [19, 50] sourced real-world underwater images and categorized them according to the degree of degradation. Unsupervised learning methods [24, 36] were developed on top of these data by learning the mapping from severely degraded images to slightly degraded images. However, these approaches were difficult to train and tended to produce unstable results. Recently, semi-supervised learning methods [38, 28, 29, 30] have been developed, using both labeled and unlabeled underwater images. Although promising results were derived, these methods did not consider physical principles.
@@ III Methodology
Considering the importance of physical principles to the UIE task, we propose a Physics-Aware Triple-Stream Underwater Image Enhancement Network (PATS-UIENet). This network contains a Direct Signal Transmission Estimation Steam (D-Stream), a Backscatter Signal Transmission Estimation Steam (B-Stream) and an Ambient Light Estimation Stream (A-Stream), which are used to explicitly estimate the three parameters of the revised IFM [2], respectively. To address the challenge of the lack of labeled real-world underwater images, we further adopt an IFM-inspired semi-supervised learning framework, which consists of a bi-directional supervised scheme and an unsupervised scheme. Compared to the supervised or unsupervised method, the PATS-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model trained is thus stronger, due to the complementary action of the two schemes.
III-A Physics-Inspired Design
Motivated by the revised IFM [2], we propose a physics-aware underwater image enhancement network on top of this model. Since the network explicitly estimates the parameters of the revised IFM by learning from training images, it avoids the rigid constraints of priors while improving the interpretability of the model trained. To reduce the complexity of the network, we simplify Eq. (2) as follows:
where () is the degraded image, is the underlying clean image, and represent the direct and backscatter signal transmission maps, respectively, and denotes the ambient light. To estimate the three parameters, including , and , we design a D-Stream, a B-Stream and an A-Stream, respectively.
We also introduce an IFM-inspired semi-supervised learning framework by leveraging the explicitly estimated parameters of the revised IFM [2]. This framework consists of a bi-directional supervised learning scheme and an unsupervised learning scheme, which addresses the challenge of rare labeled real-world underwater images. Compared to the supervised or unsupervised methods, our PATS-UIENet can take advantage of both labeled and unlabeled real-world images for the more effective training operation. Furthermore, the complementary action between the two schemes enhances the generalization ability of the model trained.
@@ III-B PATS-UIENet
Red Channel Tuner (RCT). Inspired by the channel attention mechanism [52], we propose a compact RCT module in order to adaptively emphasize the information contained in the red channel. This module first uses a convolutional layer to extract basic features from the input image, and then use the Global Average Pooling (GAP) to obtain a global representation. A fully-connected layer and the Sigmoid activation function are further used to transform this representation to a tuning weight. The weight is finally used to scale the red channel of the image.
@@
D-Stream. D-stream is designed to estimate the direct signal transmission map of an underwater image. According to Eq. (3), the transmission map varies at the locations of different pixels. In this case, the network should be able to generate fine-grained representations. Hence, we build the D-Stream using a CNN-based encoder-decoder network, due to its strong ability to learn local characteristics, which helps preserve edges and fine details. The encoder comprises five consecutive blocks, denoted as to . As a result, multi-scale feature maps are derived using the encoder. Given is used as the bottleneck, the decoder contains four blocks, denoted as to , symmetrically. Skip connection is used to pass the feature maps produced by an encoder block to the corresponding decoder block at the same level, which is useful for restoring the fine-grained spatial structure. The use of multi-scale features and skip connections enhances the ability of the network to restore the spatial structure of the scene. The output of the last decoder block is fed into a convolutional layer. Three transmission maps are produced in terms of different color channels. In essence, D-Stream is specifically focused on modeling the directly transmitted light, which normally contains the spatial structure of the scene.
B-Stream. B-stream is responsible for modeling the degradation caused by the backscattering in underwater environments. Although it shares the same CNN-based architecture as D-Stream for training stability, B-stream differs in functionality and processing strategy. Unlike direct signals, backscatter signals exhibit strong global interference and require a context-aware estimation. To this end, we pass the output of the first encoder block through a Patchify processing and feed it into the A-Stream, leveraging the global modeling capability of Transformer to enhance the estimation accuracy. Moreover, B-Stream is connected with the A-Stream via Residual Communication Modules (RCMs), enabling bidirectional feature exchange. Compared to the D-Stream, B-Stream is focused on modeling the global degradation rather than recovering the local scene content.
A-Stream. Ambient light reflects the overall luminance in the underwater scene and is generally spatially homogeneous. To effectively model this global characteristic, we employ a Transformer-based design instead of CNNs, because Transformer [53] is well-suited for capturing long-range dependencies and global characteristics through the self-attention mechanism. A-Stream consists of five Transformer blocks, denoted as to , and includes a dedicated “Ambient” token to specifically learn the ambient light parameter. Additionally, A-Stream is coupled with the B-Stream through the RCM, allowing it to incorporate local features and mitigate the lack of spatial detail inherent in Transformer. A-Stream is distinguished from the other streams by its global feature learning process and Transformer-based architecture, which complements the local characteristic modeling of the CNN-based streams.
@@ III-C IFM-Inspired Semi-supervised Learning Framework
To overcome the challenge of insufficient real-world underwater images, we adopt a semi-supervised learning framework (see Fig. 3), inspired by the revised IFM [2] theory, embedding physical constraints directly into the learning process to effectively leverage unlabeled data. This framework includes a bi-directional supervised scheme and an unsupervised scheme. The bi-directional scheme contains a forward-enhancement module, which learns from reference images and uses the estimated parameters of the revised IFM to perform UIE, and a backward-degradation module, which uses the estimated parameters to degrade the reference images to their real-world degraded counterparts. As a result, the bi-directional supervised scheme can learn a more accurate estimation of the parameters using limited real-world data than that learned using a supervised method with a large number of synthetic data. Also, the generalization of the model trained is stronger than that trained using a supervised method with limited real-world data. On the other hand, the unsupervised scheme exploits the unlabeled real-world underwater images that we collect. A second set of images with different degrees of degradation are generated from these images based on the revised IFM [2]. Two sets of parameters can be estimated using both sets of images, respectively. Each set is utilized as the reference for the other set because their IFMs are related.
@@
Unsupervised Learning. To leverage unlabeled underwater images, we adopt an unsupervised learning scheme based on the physical modeling capability of the revised IFM [2] rather than using generative networks [34, 19], which are difficult to train and often produce unstable results. This scheme adaptively constructs pairs of images with different levels of degradation from the unlabeled data.
@@
Similar to Eq. (6), the loss function for the ambient light estimated by the A-stream is defined as:
@@ IV-A Baselines
We compared the proposed method with two prior-based approaches which were developed on top of the IFM [1], including UDCP [4] and DHL [10], three prior-based methods which did not consider the IFM, including Histogram Equalization (HE) [9], HLRP [15] and MMLE [14]. We also compared our method with seven supervised learning approaches, including WaterNet [22], UColor [23], FUnIE-GAN [20], PUIE-Net (MP) [25], U-Transformer [21], URanker [26] and CCMSRNet [27], an unsupervised learning method, i.e., USUIR [24], and three semi-supervised learning methods, including DDFormer [29], Semi-UIR [28] and UWFormer [30].
@@ IV-C Evaluation Metrics
For the labeled testing data, we used Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [57] and Learned Perceptual Image Patch Similarity (LPIPS) [58] to assess the quality of an enhanced image with regard to the reference image. Regarding the unlabeled testing data, we used the Underwater Image Fidelity (UIF) [59] and Multi-scale Image Quality (MUSIQ) [60] metrics. When the SQUID [10] data set was tested, the Average Reproduction Angular Error () [10] was used to evaluate the quality of color restoration. Since transmission maps are related to the scene distance, we used the Pearson Correlation Coefficient (PCC) calculated between the estimated transmission maps and the ground-truth depth map to assess the performance of transmission estimation [10].
@@ IV-D Implementation Details
We implemented the proposed method using Pytorch and conducted experiments on Ubuntu 20.04 with a GeForce RTX 3090 graphics processing unit. During the training process, the images were resized to a resolution of pixels. The number of filters in the RCT was set to 16. For the to in the D-Stream and B-Stream, the numbers of filters were set to 64, 128, 256, 512, 512, 256, 128, 64 and 64 in turn. Regarding the A-Stream, the dimension of each token was 384 and the number of attention heads in the MHSA was 6. We trained the PATS-UIENet using the AdamW [61] optimizer. The learning rate and batch size were set to 1e-4 and 12, respectively. Four weighting factors, including and , were set to 0.1, 0.005, 1 and 0.1, respectively. We first trained the PATS-UIENet using the bi-directional supervised scheme for 50 epochs as a warm-up stage. Then the semi-supervised scheme was used to train it for 1000 epochs. When the semi-supervised learning process was carried out, in particular, the unsupervised scheme was performed for 30 iterations after the bi-directional supervised scheme was conducted for each epoch.